Operations Report – November 3rd

MDRS Operations Report 3Nov2021

Name of person filing report: Atila Meszaros

Reason for Report: Daily

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: Nothing to report

ScienceDome Dual Split: Off

Solar—Nominal, supplying all power

Notes on solar: Check supplemental report

Notes on power system: Generator didn’t run last night as expected. We are going to run it manually for the next couple of nights to assess the situation.

Ethanol Free Gasoline – 1 gallon

Water (loft tank): 50 gallons

Water Meter: 0152994.2

Water (static tank) – 240 gallons

Static to Loft Pump used – yes

Water in GreenHab – 70 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 90 gallons

Hab toilet tank emptied: No

Perseverance rover used: Yes

Hours: 223.7

Beginning charge: 100

Ending charge: 55

Currently charging: No

Sojourner rover used: Yes


Beginning charge: 100

Ending charge:

Currently charging: No

Spirit rover used: Yes

Hours: 143.7

Beginning charge: 100

Ending charge: 76

Currently charging: Yes

Opportunity rover used: Yes

Hours: 75.0

Beginning charge: 93

Ending charge: 65

Currently charging: Yes

Curiosity rover used: Yes

Hours: 151.9

Beginning charge: 100

Ending charge: 78

Currently charging: Yes

Notes on rovers: Nothing to report.

General notes and comments:

Summary of internet: All accounts are nominal.

Summary of suits and radios: One of the headsets stop operating during the EVA preparation and was replaced on time.

Summary of general operations: The one-EVA had to be split into two EVAs due to a logistic error, but the situation was rapidly solved with the support of Dr. Rupert.

Summary of Hab operations: Water alarm on the loft tank installed to prevent overflowing. Geology EVAs and debriefing through the day.

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Geology class debriefing in the afternoon after our two last EVAs and rocket building.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: The heating part for the observatory arrived.

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Nothing at this time.

Copyright © The Mars Society. All rights reserved. | Main Site